OpenStudy (anonymous):

If f(x) = 5x + 2 and g(x) = 6x - 3, what is f(g(x))?

OpenStudy (diyadiya):

Replace x with the value of g(x) f(g(x))= 5(6x-3) + 2

OpenStudy (diyadiya):

Now distribute 5 in 5(6x - 3)

OpenStudy (diyadiya):

can you try it?

OpenStudy (anonymous):

When you say f(g(x)), it means take the value of g(x) and that will equal x in f(x). To make it easier, just rewrite f(x) as 5c+2, just so it's easier to read. So: f(c) = 5c+2 g(x) = 6x-3 When you have f(g(x)), c = 6x-3 Therefore: f(c) = 5(6x-3)+2

OpenStudy (anonymous):

(30x-15)+2

OpenStudy (diyadiya):

Yes! 30x - 15 + 2 what is -15 + 2?

OpenStudy (anonymous):

-__- lol -13

OpenStudy (diyadiya):

Yes! so final answer is? ;D

OpenStudy (anonymous):

30x-13

OpenStudy (diyadiya):

Very Good! :)